OpenStudy (anonymous):

Consider the following linear system. For what values of “K “ if any, will system has (a) no solution, (b) a unique solution ? X1 + x2 + kx3 = 1 X1 + kx2 + x3 =1 Kx1 + x2 + x3 = -2

ganeshie8 (ganeshie8):

\[ \left[ \begin{array}{cccc} 1 & 1 & k & 1 \\ 1 & k & 1 & 1 \\ k & 1 & 1 & -2 \\ \end{array} \right] \]

ganeshie8 (ganeshie8):

familiar wid row reduction right ?

OpenStudy (anonymous):

right, but could not figure it out

ganeshie8 (ganeshie8):

lets reduce it and see wat we get

ganeshie8 (ganeshie8):

\[ \left[ \begin{array}{cccc} 1 & 1 & k & 1 \\ 1 & k & 1 & 1 \\ k & 1 & 1 & -2 \\ \end{array} \right] \] R2-R1 R3-kR1 \[ \left[ \begin{array}{cccc} 1 & 1 & k & 1 \\ 0 & k-1 & 1-k & 0 \\ 0 & 1-k & 1-k^2 & -2-k \\ \end{array} \right] \] R3+R2 R2/(k-1) \[ \left[ \begin{array}{cccc} 1 & 1 & k & 1 \\ 0 & 1 & -1 & 0 \\ 0 & 0 & 2-k-k^2 & -2-k \\ \end{array} \right] \]

ganeshie8 (ganeshie8):

does it look okay ?

OpenStudy (anonymous):

ok, and then ?

ganeshie8 (ganeshie8):

next is easy, lets work infinite solutions first :- if last row has all 0, then the system will have infinite solutions.

ganeshie8 (ganeshie8):

so, set the two elements to 0 and solve for k \(2-k-k^2 = 0\) \(k = ?\) \(-2-k = 0\) \(k = ?\)

ganeshie8 (ganeshie8):

\(2-k-k^2 = 0 \) \(k = -2, 1\) so this element at 33 becomes 0 when \( k = -2~~or~~ k = 1\)

ganeshie8 (ganeshie8):

\(-2-k = 0 \) \(k = -2\) so this element at 34 becomes 0 when \( k = -2 ~~ only\)

ganeshie8 (ganeshie8):

so the system will have :- 1) infinite solutions when k=-2 2) NO solutions when k = 1 3) unique solution all other cases

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

It does, tx for help

ganeshie8 (ganeshie8):

np :)

ganeshie8 (ganeshie8):

if u are familiar with rank concept, there is a nice explanation for wat we did...